Consider the given matrix


B = 2 2 0

1 0 1

0 1 1


Find detB and use it to determine whether or not B is invertible, and if so, find B-1 . (Hint: Use the matrix equation BX = I)





1
Expert's answer
2021-05-12T16:02:32-0400

det(B)=


"\begin{vmatrix}
 2& 2 & 0\\
 1 & 0 & 1 \\
0 & 1 & 1
\end{vmatrix}" = -2 -2 = -4 "\ne0" means inverse exists


 BX = I

x = B-1I


x= "\left[\begin{array}{ccc|ccc}2 & 2 & 0 & 1 & 0 & 0\\1 & 0 & 1 & 0 & 1 & 0\\0 & 1 & 1 & 0 & 0 & 1\end{array}\right]"

Divide row 1 by 2 ": R_{1} = \frac{R_{1}}{2}"


"\left[\begin{array}{ccc|ccc}1 & 1 & 0 & \frac{1}{2} & 0 & 0\\1 & 0 & 1 & 0 & 1 & 0\\0 & 1 & 1 & 0 & 0 & 1\end{array}\right]"


Subtract row 1 from row 2 :"R_{2} = R_{2} - R_{1}"


"\left[\begin{array}{ccc|ccc}1 & 1 & 0 & \frac{1}{2} & 0 & 0\\0 & -1 & 1 & - \frac{1}{2} & 1 & 0\\0 & 1 & 1 & 0 & 0 & 1\end{array}\right]"


"R_{2} = - R_{2}"


"\left[\begin{array}{ccc|ccc}1 & 1 & 0 & \frac{1}{2} & 0 & 0\\0 & 1 & -1 & \frac{1}{2} & -1 & 0\\0 & 1 & 1 & 0 & 0 & 1\end{array}\right]"


"R_{1} = R_{1} - R_{2}"

"\left[\begin{array}{ccc|ccc}1 & 0 & 1 & 0 & 1 & 0\\0 & 1 & -1 & \frac{1}{2} & -1 & 0\\0 & 1 & 1 & 0 & 0 & 1\end{array}\right]"


"R_{3} = R_{3} - R_{2}"

"\left[\begin{array}{ccc|ccc}1 & 0 & 1 & 0 & 1 & 0\\0 & 1 & -1 & \frac{1}{2} & -1 & 0\\0 & 0 & 2 & - \frac{1}{2} & 1 & 1\end{array}\right]"



"R_{3} = \frac{R_{3}}{2}"

"\left[\begin{array}{ccc|ccc}1 & 0 & 1 & 0 & 1 & 0\\0 & 1 & -1 & \frac{1}{2} & -1 & 0\\0 & 0 & 1 & - \frac{1}{4} & \frac{1}{2} & \frac{1}{2}\end{array}\right]"


"R_{1} = R_{1} - R_{3}"

"\left[\begin{array}{ccc|ccc}1 & 0 & 0 & \frac{1}{4} & \frac{1}{2} & - \frac{1}{2}\\0 & 1 & -1 & \frac{1}{2} & -1 & 0\\0 & 0 & 1 & - \frac{1}{4} & \frac{1}{2} & \frac{1}{2}\end{array}\right]"



"R_{2} = R_{2} + R_{3}"

"\left[\begin{array}{ccc|ccc}1 & 0 & 0 & \frac{1}{4} & \frac{1}{2} & - \frac{1}{2}\\0 & 1 & 0 & \frac{1}{4} & - \frac{1}{2} & \frac{1}{2}\\0 & 0 & 1 & - \frac{1}{4} & \frac{1}{2} & \frac{1}{2}\end{array}\right]"



"ANSWER" =

"\left[\begin{array}{ccc}\frac{1}{4} & \frac{1}{2} & - \frac{1}{2}\\\frac{1}{4} & - \frac{1}{2} & \frac{1}{2}\\- \frac{1}{4} & \frac{1}{2} & \frac{1}{2}\end{array}\right] \\ = \left[\begin{array}{ccc}0.25 & 0.5 & -0.5\\0.25 & -0.5 & 0.5\\-0.25 & 0.5 & 0.5\end{array}\right]" = "B^{-1}"